OpenStudy (anonymous):

please help trying to find the end points for the circle. (x-2)^2+(y-6)^2=16

ganeshie8 (ganeshie8):

you will never make it, a circle has no end points..

ganeshie8 (ganeshie8):

could you tell what you mean by endpoints ?

OpenStudy (anonymous):

sorry i meant when x and y pass 0. x=0 and y =0 so (0, ) ( ,0)

ganeshie8 (ganeshie8):

plugin x=0 in the equation and solve y

ganeshie8 (ganeshie8):

(x-2)^2+(y-6)^2=16 (0-2)^2+(y-6)^2=16

ganeshie8 (ganeshie8):

can you solve y ?

OpenStudy (anonymous):

so i just plug in y? i dont have to divide by 16?

ganeshie8 (ganeshie8):

do you have options ?

OpenStudy (anonymous):

y=+-4.47?

ganeshie8 (ganeshie8):

nope, im getting y = 2.54, 9,46

ganeshie8 (ganeshie8):

and there are no x intercepts.. can you check the equation again ?

OpenStudy (anonymous):

sorry i didnt type it (x-2)^2+(y-6)^2=16

ganeshie8 (ganeshie8):

could you take a screenshot of the complete question and attach ?

ganeshie8 (ganeshie8):

??